OpenStudy (anonymous):

The gas tank of a truck is a cylinder 4.5ft long with a diameter of 2.5ft . At the gas station, a pump pours gas at the rate of 5ft^3 per minute. How many minutes does it take to fill the empty tank with that pump? Use the value 3.14 for Pi , and round your answer to the nearest minute.

OpenStudy (anonymous):

You need to work out the volume of the Cyclinder so it will be \[\pi r^2 h\]\[= 3.14 \times (2.5/2)^2 \times 4.5\] = 22.078125 ft^3 22.078125/5.000000 = 4.415625 minutes